OpenStudy (anonymous):

What number must you add to the polynomial below to complete the square? x2 - 3x

OpenStudy (fantageplayer):

kid buu?

OpenStudy (fantageplayer):

@2kende

OpenStudy (mathstudent55):

If you square the binomial x - a, you get: \((x - a)^2 = (x - a)(x - a) = x^2 - 2a + a^2\)

OpenStudy (mathstudent55):

Compare what you have with the square of the binomial.